Greenhab Report – April 6th

GreenHab Officer: Cheyenne Chamart

Environmental control :

Average temperatures: 78°F
Hours of supplemental light: N/A
Daily water usage for crops: N/A
Daily water usage for research and/or other purposes: N/A
Water in Blue Tank (full – not used) gallons
Time(s) of watering for crops: N/A

Changes to crops: N/A

Narrative: I stayed and worked all day in the GreenHab today with some music. I let the door open to get some fresh air, so it was very breathable and less hot. I checked the parameters and measured my plants as always. Everything is going fine, nothing to indicate!

Harvest: N/A

Support/supplies needed: Nothing.

Research Report – April 6th

Astronomy Report

Name: Ignacio Sánchez Casla
Crew: 227

Sky Conditions: Clear
Wind Conditions: windy

Observation Start Time: None
Observation End Time: None

Summary: Solar Observatory was not used today

Objects Viewed: None
Problems Encountered: None


Robotic Telescope Requested: MLC-RCOS16

Objects to be imaged this evening:
– Messier 51
– Messier 64
– Messier 57
– Messier 101
– NGC 7023
same as always
Images submitted with this report: None, still working on the Photoshop

Errors Encountered: None

Thank you for the tips that you gave me yesterday. I am going with the Photoshop option although I am still working on it. I may have been over-processing the images…

Journalist Report – April 6th

[Clear Sky] Journalist Report
Crew 227 Journalist Report 6 April 2022
Sol : 10
Name of the person filing the report: Sirga Drouet

Hello Earth, this is Mars on sol 10.

Imagine, putting on your scarf, and going out of a lighted room to the Observatory.

The moment when your eyes acclimatize to the brightness is incredible. In the beginning, a few lights, then the dark night turns into a Milky Way where the number of luminous points is incalculable, the shooting stars pass by, we are there to contemplate the universe which offers itself to us, suggests us a tiny part of its immensity. This vision gives us the desire to travel through space and not just stand there.

We are there to cultivate the silence without anything interfering and time is no longer a dimension in our eyes. This is the moment we experienced last night before giving way to sleep.

This morning the dawn came to tickle our sleeping eyes and the feeling that everything is possible in the first hours of the day takes us back.

But we have other projects that take us out of our dreams.

Today EVA walked to North Bridge for Julien’s last game against me. We climbed hills that changed color as the day goes by. This morning it was red then orange. As we climbed, the sun brightened them more and more and we were at the challenge point.

Thanks to our training we were faster, more efficient and confident. We went straight back home after this escapade. After lunch, everyone had their stuff. For my part I made cookies with my accomplice for the snack without saying that they were filled with insects, the team saw only fire and my stratagem worked wonderfully hehe.

The day passed at one of its speeds and it is already time for Cap Comm.

It’s always a pleasure to wake up in March and I hope that tomorrow will be fruitful.

See you soon

Supplemental Operations Report – April 6th

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Power system monitor

Action taken for non-nominal systems: Still figuring out the best option for it. It needs to be hardwired into the internet, but also close enough to the ScienceDome to connect to the data

Generator: Still being manually run at night.

ScienceDome Dual Split: Off

Solar—Charge controllers off at night.

Solar— Nominal

Notes on power system: The power system is acting crazy again. Yesterday we lost power due to another overload of the solar system. Then today the SOC is reading incorrectly again.I will look at it between crews.

Propane Reading, station tank – 60%

Propane Reading, director tank— 60%

Propane Reading, intern tank— 70%

Propane Reading, generator— 32 %

Ethanol Free Gasoline – 0 gallons.

Water (Outpost tank) – 250 gallons

Sojourner rover used: no

Hours: 167.5

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: Still haven’t started on the battery cleaning and watering.

ATV’s Used: (Honda, 350.1, 350.2, 300): none

Reason for use: n/a

Oil Added? No.

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used: 0 hours

Notes on ATVs: 350.1 is blocking the gap between the campus and the outpost so no one will damage the repaired wiring conduit.

HabCar used and why, where? Yes, to town for garbage and water runs

CrewCar used and why, where? Yes, to town

Luna used and why, where? Yes, to town

Campus-wide inspection, if action is taken, what and why? Nothing to report.

General notes and comments: Nothing to report

Summary of general operations: Nothing to report

Summary of internet: We removed the modems and routers for the two HughesNet systems that have been deactivated in the Hab. Starlink will now be the only internet available on campus for crews. We do need to secure it with a password still.

Summary of suits and radios: Nothing to report

Summary of Hab operations: Evidence of rodents noted.

Summary of Outpost operations: Evidence of rodents noted.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Went into the robotic observatory today and restarted the system with Peter. Also realigned the small telescope’s camera which had become loose where the cords connect.

Summary of health and safety issues: Nothing to report.

Commander Report – April 6th

Crew 227 Commander Report 6th April 2022
SOL: 10
Name of person filing report: Cyril Wain

Cookie day today!

The day started with breakfast at 08h00 with the breakfast. We ate some cereals with dried milk and dried apples with some cinnamon. As usual, Julien studied our sleep and Jean did some analysis of our metabolism.

Last night we caught a huge rodent so, during breakfast, we had a discussion to know what to do with him: kill him or release it during the EVA. The concern was that if we released him we were afraid he would end up coming back but on the other hand half of the team didn’t want to kill him. So, we decided to release him and we hope he won’t come back.

At 9h00, a part of the crew went on EVA. The goal was to perform a match on the experiment of Julie. This time it was Sirga against Julien. Therefore, the group was constituted of Julien, Sirga, Jean, Julie, and Audrey. Julien won the match. The EVA ends at 10h30. Julie and Sirga have extended the EVA in order to release the rodent. The crew was happy that this rodent is now out of the station.

At lunch, we ate some soup and then some fried potatoes. After, everyone worked on his/her personal experiment.

At 16h00, Julien and Sirga cooked some cookies and Sirga added her best ingredient: some worms! In the end, we don’t really taste it and the cookies were really tasty!

Then, we worked again on our different experiments until 18h30.

Now, we are doing our reports for the Capom window.

Tonight, we would like to continue our sky observations and try to take some time-lapse of the stars.

Regarding the mood of the team, it was a bit tense at the breakfast with the rodent argument but everything quickly went down and with the cookies, the morale of the team is now as good as usual.

Regarding the station, except for the invasion of rodents (we can still hear other ones) and a small issue with the SOC (now fixed), everything is nominal.

Operations Report – April 6th

Crew 227 Operations Report 06-04-2022

SOL: 10

Name of person filing report: Julien Meert

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report


Spirit rover used: Not use

Opportunity rover used: Yes

Hours: 93,1

Beginning charge: 100%

Ending charge: 77

Currently charging: plugged in

Curiosity rover used: Not used

Perseverance rover used: Not used

General notes and comments: Nothing to report

Summary of Hab operations : We had some rodents in the Hab and yesterday night was quite noisy, so we put out our forks and went to war ! Armed with an empty pot and some cheese as a trap we try to ambush them around midnight!

When we left the Upper part to go to the bathroom, the previous plan perfectly worked ! We caught our invisible guest ! (Which was not that invisible in order to the size of the plenty (of a piece of wall maybe) belly that the enemy got !)

After a big debate at the breakfast, about the way that we are going to get rid of our night prisoners, we finally decide with Shannon that we are going to make him do the walk of shame in the middle of the desert.


Water (static tank): 276 gallons

Water (loft tank): 25 gallons

Water Meter: 157308,1 units

Static to Loft Pump used : 0 gallons

Static tank pipe heater (on or off): Off

Static tank heater Off

Toilet tank emptied: Yes

Summary of internet: No problems today

Summary of suits and radios: Nothing new

Summary of GreenHab operations: Nothing new

WATER USE: 0 gallons

Heater: Off

Supplemental light: Off

Harvest: 0

Summary of ScienceDome operations: Audrey used the oden today.

Dual split: Off

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support : Nothing special for today !

Copyright © The Mars Society. All rights reserved. | Main Site